what is the slope if the line is -1/2 and it passes through the point (1,2)?
Answer by plourde216(7)

When you are given the slope of a line and a point that is on the line, you should use point-slope form. Point-slope form is (y-y1)=m(x-x1). The m stands for slope, x1 and y1 are the values from the ordered pair that you are give. In this case, the formula would read (y-2)=(-1/2)(x-1). Your next step would be to distribute the -1/2 to (x-1). This will leave you with (-1/2)x + (1/2) on the right side of the equation.
y-2=(-1/2)x + (1/2)
Now you will add 2 to both sides in order to get y by itself.
y=(-1/2)x +(5/2)
Hope this helps